OpenStudy (please):

Can you help me solve this problem???? How many carbon atoms are there in 2.12 molecules of propane? (C3H8)

OpenStudy (jfraser):

Since you can't have 2.12 MOLECULES of propane, I assume you mean 2.12 MOLES of propane. Since there are always 6.022*10^23 molecules in 1 mole of molecules, 2.12 moles will have \[2.12 moles * (\frac{6.022*10^{23} molecules}{1 mole}) = 1.28*10^{24} molecules\] in that many moles of propane. Since every molecule of prpane contains 3 atoms of carbon, the total number of carbon atoms in 2.12 moles of propane is\[12 moles * (\frac{6.022*10^{23} molecules}{1 mole})*(\frac{3 C atoms}{1 C_3H_* molecule}) = 3.83*10^{24} C atoms\]

OpenStudy (please):

Where did you get the 12 moles from?

OpenStudy (please):

@JFraser

OpenStudy (anonymous):

it's not 12 it's 2,12...2 is not seen..

OpenStudy (jfraser):

^right, I type too fast. It should be 2.12 moles, not 12

OpenStudy (anonymous):

yes i thought the same..answer is correct)